# Exploring a tentative link between FRBs and pulsars with broad energy distributions and applications for nearby FRB survey strategies

### v1

@@ Abstract
Fast radio bursts (FRBs) are energetic, short-duration radio pulses of unclear origins. In this study, we investigate the FRBs and pulsars with broad energy distributions by fitting their high energy tails with a power-law model. Two cosmological repeating FRBs (FRB 20201124A and FRB 20220912A), one nearby FRB (FRB 20200120E), and two pulsars (RRATs J1846--0257 and J1854+0306), exhibit power-law indices of $\alpha\gtrsim-1$, suggesting that their bright pulses contribute significantly to the total radio pulse energy. The brightest bursts from these sources fit well with a simple power-law model ($\alpha=-0.26\pm 0.05$), indicating a tentative link between certain high-luminosity FRBs and low-luminosity radio bursts. We also discuss detailed survey strategies for FAST, MeerKAT and Parkes cryoPAF in the search for FRBs in nearby globular clusters (GCs) using different power-law indices, recommending targets for observation. We suggest that combining observations with FAST ($\sim 3$ hours) and Parkes cryoPAF (10–20 hours) are practicable for discovering new FRBs in the nearby GCs.
@@ 1 Introduction
Fast radio bursts (FRBs) are immensely energetic radio pulses with durations ranging from microseconds to seconds (Lorimer et al., 2007; Snelders et al., 2023; Chime/Frb Collaboration et al., 2022). Among the nearly one thousand known FRB sources, fewer than 10% exhibit repeat bursts (CHIME/FRB Collaboration et al., 2021; Spitler et al., 2016; Chime/Frb Collaboration et al., 2023). The origin of FRBs remains unclear, but the only two nearby events offer clues to their engines: the apparently one-off burst FRB 20200428 from the Galactic magnetar SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020) and the repeating FRB 20200120E, localized to a globular cluster (GC) in the M81 galactic system at a distance of 3.63 Mpc (Bhardwaj et al., 2021; Kirsten et al., 2022). While these two events suggest connections to distinct environments, young magnetars and ancient star populations, respectively, both have propelled FRB theories involving compact objects, especially neutron stars, as potential sources (Zhang, 2023).
@@
The typical energy of cosmological repeating FRBs (Li et al., 2021; Xu et al., 2022; Chime/Frb Collaboration et al., 2020) is approximately four orders of magnitude greater than that of the two nearby FRBs (Bochenek et al., 2020; Nimmo et al., 2023), and about twelve orders of magnitude larger than that of pulsars [^1] http://www.atnf.csiro.au/research/pulsar/psrcat (Manchester et al., 2005). However, the detection of a bright burst from FRB 20200120E (Zhang et al., 2024a) bridges the gap between nearby and cosmological FRBs. Similarly, the detection of radio bursts with varying luminosities from SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Kirsten et al., 2021; Good & Chime/Frb Collaboration, 2020; Zhu et al., 2023) connects the energy range of nearby FRBs to that of pulsars. As presented in Figure 1 of Zhang et al. (2024a), no clear gap exists between the estimated isotropic-equivalent energy releases of various radio pulses.
It is intriguing to explore whether nearby radio pulse phenomena can be connected to cosmological FRBs, particularly repeating ones. Follow-up observation of repeating FRBs (Li et al., 2021; Xu et al., 2022; Zhang et al., 2023) have revealed extremely variable properties, especially in energy, spanning more than three orders of magnitude. In contract, broad energy distributions are atypically for single pulses from pulsars (Burke-Spolaor et al., 2012). Intensity variations exceeding an order of magnitude are usually attributed to the rare phenomenon of “giant pulses” (Kuzmin, 2007), reported in only about ten (Wang et al., 2019) of the more than 3500 known pulsars (Manchester et al., 2005). However, recent observations of some of another kind of special pulsars (Zhang et al., 2024b, 2024), rotating radio transients (RRATs), characterized by sporadic radio pulses, have also shown broad energy distributions. It is still worthy to explore the potential relation between FRBs and these pulsars with broad energy distributions.
A few hyperactive repeating FRBs provide large samples of $\gtrsim 1000$ burst, enabling detailed studies of their energy distributions (Li et al., 2021; Xu et al., 2022; Zhang et al., 2023). Notably, two such sources–FRB 20201124A and FRB 20220912A–have been observed for more than 2000 hours, revealing breaks in their burst energy distributions, with a flattening of the power-law slope ($\gtrsim-1$) at higher energies (Kirsten et al., 2024; Ould-Boukattine et al., 2024). It is also noteworthy that the detection of a bright burst from the nearby FRB 20200120E indicates a similar flat energy distribution, with a power-law index of approximately -0.98 (Zhang et al., 2024a). Furthermore, two RRATs have also exhibited power-law-like tails at high energies with similarly flat indices (Zhang et al., 2024). These observations suggest that flat power-law tails at high energies may offer a clue into the mechanisms underlying these short-duration radio pulses with diverse energy releases.
In this study, we present an analysis of the energy distributions tails at high energy of FRBs and pulsars using a consistent method. Section 2 describes the sources selection and data processing, while Section 3 and 4 presents and discusses the results. We conclude in Section 5.
2 Sources selection and data processing
2.1 Sources selection
Figure 1: Isotropic-equivalent spectral luminosity versus the product of observing frequency and pulse width for various pulsars and FRBs. The brightest strong pulses from RRATs with relatively high $B_{\rm s}$ (J1846-0257 and 1854+0306) are represented by red and black filled stars, while their weak emissions are denoted by unfilled stars (Zhang et al., 2024). The brightest giant pulses from J0534+2200 (Crab pulsar) (Bera & Chengalur, 2019) and J0540--6919  (Geyer et al., 2021) are shown as blue and violet filled triangles, with their normal emissions  (Manchester et al., 2005) marked by unfilled triangles. The brightest burst from the nearby FRB 20200120E (Zhang et al., 2024a) is indicated by an orange filled square, with its relatively weaker bursts (Nimmo et al., 2023) are marked by unfilled squares. The brightest bursts from cosmological FRB 20201124A (Kirsten et al., 2024) and FRB 20220912A (Ould-Boukattine et al., 2024) are represented by black and gray filled circles, with their relatively weaker bursts (Xu et al., 2022; Zhang et al., 2023) shown as unfilled circles. The pulsar (Manchester et al., 2005) population is indicated in thistle. Grey dashed lines represent constant isotropic-equivalent energy release, while grey dotted lines indicate constant brightness temperature.
The selected sources, presented in Figure 1, show the brightest strong pulses that are at least 2--3 orders of magnitude more energetic than their relatively weak emissions. The energies of the detected radio pulses from these sources span roughly 15 orders of magnitude. Specifically, the studied sources include: (1) FRB 20201124A and FRB 20220912A: These are the only two FRBs with up to 2000 hours of follow-up observations, during which ultra-luminous bursts with isotropic energy up to 1041 erg were detected (Kirsten et al., 2024; Ould-Boukattine et al., 2024), placing them among the most energetic FRBs observed. (2) FRB 20200120E: A nearby FRB that exhibited a burst storm (Nimmo et al., 2023) and an abnormal burst more than 42 times brighter than previously detected bursts at a similar frequency (Zhang et al., 2024a). (3) J0534+2200 (the Crab pulsar) and J0540--6919: These pulsars are known for producing giant pulses with the largest excesses relative to their normal emissions (Bera & Chengalur, 2019; Geyer et al., 2021; Wang et al., 2019). (4) RRATs J1846--0257 and J1854+0306: These RRATs have relatively high surface magnetic fields ($\sim 2\times 10^{13}$ G) and show weak emissions that are up to 687 and 512 times fainter than their strong individual pulses. We also analysed data from RRAT J1913+1330 (Zhang et al., 2024b), which exhibits extreme pulse-to-pulse modulation and an energy distribution spanning 3 orders of magnitude but could be well-fitted by two log-normal functions. Its distribution tail at the high energy part provides a useful comparison with the four source kinds above.
@@ 2.2 Data processing
where $D_{L}$ is the luminosity distance, $z$ is the redshift, and $F$ is the pulse fluence. The energy values for each source were then normalized to their brightest bursts.
Since only one bright burst has been detected from FRB 20200120E (Zhang et al., 2024a) and detailed burst properties for FRB 20220912A are currently unavailable (Ould-Boukattine et al., 2024), we were unable to reprocess their high-energy distributions. To compare our results with their reported distributions, which are cumulative, our initial analysis used a cumulative distribution, adjusting the bins for each source to cover a similar range in log-10 space. We also provide a differential distribution analysis. To enable comparisons between different sources, we defined a density that is calculated as the number of bursts per bin divided by the total number of detected bursts.
The power-law-like tails were fitted using the Python package curve_fit function from the SciPy library (Virtanen et al., 2020). For data with complex uncertainties, we performed the fitting within a Bayesian framework using Markov Chain Monte Carlo (MCMC) sampling, via the emcee Python package (Foreman-Mackey et al., 2013).
@@ 3 Results
Figure 2: Cumulative specific energy distributions of high-energy bursts from FRB 20201124A, single pulses from RRATs J1913+1330, J1846--0257, and J1854+0306, and giant pulses from the Crab pulsar (J0534+2200) and J0540--6919. The energy values for each source are normalized to their respective brightest bursts. Each high-energy tail is fitted with a simple power law model.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: Differential specific energy distributions of high-energy bursts from FRB 20201124A, single pulses from RRATs J1913+1330, J1846--0257, and J1854+0306, and giant pulses from the Crab pulsar (J0534+2200) and J0540--6919. The energy values for each source are normalized to their respective brightest bursts. Each high-energy tail is fitted with a simple power law model.
Figure 4: Comparison of cumulative (top) and differential (bottom) normalized specific energy distributions of high-energy bursts from FRB 20201124A, single pulses from RRATs J1913+1330, J1846--0257, and J1854+0306, and giant pulses from the Crab pulsar (J0534+2200) and J0540--6919. The energy values for each source are normalized to their respective brightest bursts. The density represents the number of bursts per bin divided by the total number of detected bursts.
Figure 2 presents the cumulative specific energy distributions of high-energy bursts from the selected sources with relatively sufficient burst counts, most of which can be well-fitted by a simple power-law model. The high-energy tail of the distribution for RRAT J1913+1330 is forcefully fitted using a power-law function, although the fit is not as good as for the other sources, in order to enable comparison. Specifically, the fitted power-law indices ($\alpha$) for FRB 20201124A (46 bursts  [^2] For events detected simultaneously by multiple telescopes, we selected bursts from the observation with the largest bandwidth.), RRATs J1846--0257 (72 bright pulses) and J1854+0306 (325 bright pulses) are $-0.75\pm 0.14$, $-0.84\pm 0.10$ and $-0.76\pm 0.03$, respectively, all of which are $\gtrsim-1$. In contrast, the indices for the Crab pulsar ($-1.98\pm 0.07$, 1955 giant pulses), J0540--6919 ($-2.29\pm 0.07$, 865 bright single pulses) and RRAT J1913+1330 ($-3.04\pm 0.29$, 1955 individual pulses) are significantly steeper, at $\lesssim-2$. Although the available data for bursts from FRB 20200120E and FRB 20220912A are insufficient to reprocess their high-energy distributions, the reported fitted power-law indices are $-0.98\pm 0.06$ (Zhang et al., 2024a) for FRB 20200120E and range from $-0.99$ to $-0.60$ (Ould-Boukattine et al., 2024) for FRB 20220912A at around 1.4 GHz, both indicate relatively flat indices.
Furthermore, we also present the differential specific energy distributions of high-energy bursts from these source in Figure 3. Similar to the cumulative distributions, the fitted indices using a power-law function for the differential distributions at the high energy tails ($\beta$) of FRB 20201124A ($-0.40\pm 0.21$), RRATs J1846--0257 ($-0.63\pm 0.16$) and J1854+0306 ($-0.65\pm 0.05$) are significantly flatter than those of the Crab pulsar ($-1.98\pm 0.09$), J0540--6919 ($-2.0\pm 0.1$) and RRAT J1913+1330 ($-2.72\pm 0.31$). However, the sample sizes for some sources, especially RRAT J1846--0257 and FRB 20201124A, are limited, and the fitted uncertainties are relatively large.
As shown in the comparison in Figure 4, the energy distributions of FRB 20201124A, RRATs J1846--0257 and J1854+0306 decline much more slowly at high energies compared to those of the giant-pulse pulsars and RRAT J1913+1330.

### v2

# Flat tails in FRB and pulsar energy distributions: implications for optimizing nearby FRB surveys

@@ Abstract
Fast radio bursts (FRBs) are energetic, short-duration radio pulses of unclear origin. To explore effective survey strategies for detecting FRBs from nearby globular clusters (GCs), we investigate the burst energy distribution, which has a strong influence on the detection rate. We re-analyze FRBs and pulsars exhibiting broad energy distributions by fitting their high-energy tails with power-law models. Two cosmological repeating FRBs (FRB 20201124A and FRB 20220912A), one nearby FRB (FRB 20200120E), and two pulsars (RRATs J1846-0257 and J1854+0306), exhibit power-law indices of $\alpha\gtrsim-1$, suggesting that their bright pulses contribute significantly to the total radio pulse energy. The brightest bursts from these sources can be fitted with a power-law model ($\alpha_{\rm Bri}=-0.26\pm 0.05$), suggesting that an extremely flat index is required to naturally derive high-luminosity FRBs from low-luminosity sources. We present detailed survey strategies for FAST, MeerKAT and Parkes cryoPAF in the search for FRBs in nearby GCs using different power-law indices, recommending targets for observation. We suggest that combining observations with FAST ($\sim 3$ hours) and Parkes cryoPAF (10-20 hours) is practicable for discovering new FRBs in nearby GCs.
@@ 1 Introduction
Fast radio bursts (FRBs) are immensely energetic radio pulses with durations ranging from microseconds to seconds (Lorimer et al., 2007; Snelders et al., 2023; Chime/Frb Collaboration et al., 2022). Among the nearly one thousand known FRB sources, fewer than 10% exhibit repeat bursts (CHIME/FRB Collaboration et al., 2021; Spitler et al., 2016; Chime/Frb Collaboration et al., 2023). The origin of FRBs remains unclear, but the only two nearby events offer clues to their engines: the apparently one-off burst FRB 20200428 from the Galactic magnetar SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020) and the repeating FRB 20200120E, localised to a globular cluster (GC) in the M81 galactic system at a distance of 3.63 Mpc (Bhardwaj et al., 2021; Kirsten et al., 2022). While these two events suggest connections to distinct environments, young magnetars and ancient star populations, respectively, both have propelled FRB theories involving compact objects, especially neutron stars, as potential sources (Zhang, 2023).
Nearby FRBs are crucial in understanding FRBs, as they enable precise localisation to identify their surroundings and deep multiwavelength follow-up to resolve their mechanisms. Although CHIME, the current most efficient FRB hunter, has been operating since 2018 and can discover FRBs daily (CHIME/FRB Collaboration et al., 2021), only one repeating FRB with a distance $\leq$ 100 Mpc has been detected (Bhardwaj et al., 2021), indicating that a blind survey is challenging for providing a sufficient sample of nearby repeating FRBs. As the only nearby repeating FRB 20200120E has been localised to a GC (Kirsten et al., 2022), Kremer et al. (2023) presented a dedicated FRB survey of GCs in nearby galaxies, with M87 being identified as the most promising target, where high-sensitivity radio telescopes like FAST or MeerKAT have a 90% probability of detecting a globular cluster FRB within approximately 10 hours of observation.
However, it is notable that the detection rate of nearby FRBs from such dedicated surveys is significantly influenced by the burst energy distribution (Kremer et al., 2023). The nearby radio bursts expected to be detected by high-sensitivity telescopes such as FAST (Jiang et al., 2020) would have relatively low energies, even comparable to pulsar single pulses. The expected event rate from surveys such as those proposed by Kirsten et al. (2022) assumes a consistent distribution from bright to faint target sources. Therefore, it is important to investigate the relationship between bright sources, such as cosmological FRBs, and relatively faint sources, such as nearby FRBs or even bright pulsar single pulses.
FRB 20200120E stands as the sole nearby repeating FRB at a distance of approximately 3.63 Mpc that enables detailed energy distribution analysis. Observations with the Effelsberg telescope recorded 65 bursts (Nimmo et al., 2023; Kirsten et al., 2022) with fluences ranging from 0.04–0.71 Jy ms at 1.4 GHz, including a remarkable “burst storm” event of 53 bursts detected within 40 minutes. The burst storm exhibited a steep energy distribution characterized by a power-law tail with index of $\sim-2.39$ (Nimmo et al., 2023). However, a recent observation have detected an exceptionally bright burst (fluence $\approx 30$ ms at 1.1-1.7 GHz), 42 times brighter than previous detection at 1.4 GHz, suggesting a flatter high-energy tail with index of approximately 0.98-0.98 (Zhang et al., 2024b).
The typical energy of cosmological repeating FRBs (Li et al., 2021; Xu et al., 2022; Chime/Frb Collaboration et al., 2020) is approximately four orders of magnitude greater than that of FRB 20200120E (Nimmo et al., 2023). The study of energy distributions in cosmological FRBs began with FRB 20121102A, the first known repeater (Spitler et al., 2016). As summarized by Wang & Zhang (2019), the early repeating samples of FRB 20121102A showed cumulative energy distributions of power-law tails with indices from 0.6~-0.6 to 0.8-0.8111It is noteworthy that their definition of the indices differs from that in our study, we have converted their values to match our definition.. Later analysis with larger samples also indicate flatter high energy tails with indices larger than 1.0-1.0 (Cruces et al., 2021; Hewitt et al., 2022; Jahns et al., 2023; Li et al., 2021). Although monitoring of other active repeaters such as FRB 20180916B and FRB 20190520B showed steep indices $($$\alpha<-2.0$) Chime/Frb Collaboration et al., 2020; Zhang et al., 2022), recent studies of two hyperactive repeaters, FRB 20201124A and FRB 20220912A, both detected extreme bright pulses by ultra-long monitoring (>2000>2000 hours), revealing breaks in their burst energy distributions, with a flattening of the power-law slope ($\gtrsim-1$) at higher energies (Kirsten et al., 2024; Ould-Boukattine et al., 2024). These flatter indices at high energies mirror FRB 20200120E’s bright burst properties.
Furthermore, the latest rate analysis of the new detected hyperactive repeater FRB 20240114A also indicates a potential flattening at higher energies (Huang et al., 2025).
The typical energy of pulsars (Manchester et al., 2005) is about twelve orders of magnitude smaller than that of cosmological repeating FRBs (Li et al., 2021; Xu et al., 2022; Chime/Frb Collaboration et al., 2020). The broad energy distributions observed in FRBs are also atypically for single pulses from pulsars (Burke-Spolaor et al., 2012). Intensity variations exceeding an order of magnitude are usually attributed to the rare phenomenon of “giant pulses” (Kuzmin, 2007), reported in only about ten (Wang et al., 2019) of the up to 4000 known pulsars [^2] ATNF Pulsar Catalogue v2.6.1: https://www.atnf.csiro.au/research/pulsar/psrcat (Manchester et al., 2005). Pulsar giant pulses normally exhibit steep energy distributions. For instance, J0534+2200 (the Crab pulsar) and J0540-6919 are known for producing giant pulses with the largest excesses relative to their normal emissions (Bera & Chengalur, 2019; Geyer et al., 2021; Wang et al., 2019). The Crab pulsar (J0534+2200) demonstrates a single power-law distribution with $\alpha\sim-3.0$ across three decades of fluence. Similarly, J0540-6919 shows a cumulative distribution with $\alpha=-2.75\pm 0.02$. However, recent observations of another kind of special pulsars (Zhang et al., 2024c, a), rotating radio transients (RRATs), characterized by sporadic radio pulses, have also shown broad energy distributions. Although one of which (J1913+1330) present steep tails (Zhang et al., 2024c), two other RRATs (J1846-0257 and J1854+0306) have also exhibited power-law-like tails at high energies with similarly flat indices ($\gtrsim-1$; Zhang et al., 2024a). These indices match the flatter high-energy regimes of FRB 20200120E and cosmological repeaters like FRB 20201124A and FRB 20220912A.
@@
The detection of a bright burst from FRB 20200120E (Zhang et al., 2024b) bridges the gap between nearby and cosmological FRBs, and the detection of radio bursts with varying luminosities from SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Kirsten et al., 2021; Good & Chime/Frb Collaboration, 2020; Zhu et al., 2023) connects the energy range of nearby FRBs to that of pulsars. No clear gap exists between the estimated isotropic-equivalent energy releases of various radio pulses (Zhang et al., 2024b), making it natural to explore the burst energy distribution relationship across radio sources with various energy releases.
Recent studies have shown that flat tails in pulse energy distributions are observed across cosmological FRBs, the nearby FRB 20200120E, and certain pulsars. However, analysis of different sources has employed various methods: some studies derived power-law indices from cumulative distributions, while others used differential distributions. Energy calculations have varied depending on whether they incorporated central frequency, bandwidth, or burst frequency extent scaling. Additionally, the choice of distribution bins and fitting tools can affect the obtained index values. In this study, we present a consistent analysis of the high-energy tails in the energy distributions of FRBs and pulsars using a unified methodological approach. Section 2 describes the source selection and data processing, while Section 3 and 4 presents and discusses the results. We conclude in Section 5.
2 Source selection and data processing
2.1 Source selection
Our selection prioritizes sources that exhibit flat high-energy tails in their burst energy distributions. Among FRBs, we include four cosmological sources with demonstrated power-law indices of $\alpha\gtrsim-1$ at high energies: FRB 20121102A, FRB 20201124A, FRB 20220912A, and FRB 20240114A. We also include the nearby FRB 20200120E, which shows a tentative high-energy tail ($\alpha\sim-0.98$), a key target of our study. For pulsars, we focus on J1846-0257 and J1854+0306, two RRATs exhibiting flat tails, in contrast to the steep energy distributions typically observed in giant pulse pulsars. For comparison, we selected the well-studied giant pulse pulsars J0534+2200 (the Crab pulsar) and J0540-6919, known for producing giant pulses with the largest excesses relative to their normal emissions. However, detailed burst properties for FRB 20220912A and FRB 20240114A are currently unavailable, limiting our ability to conduct comprehensive energy distribution analyses. Similarly, FRB 20200120E has only one documented ultra-bright burst, which restricts statistical analysis. Therefore, our quantitative energy distribution analysis focuses on the data of FRB 20121102A (Li et al., 2021, ; the largest available sample) and FRB 20201124A (Kirsten et al., 2024, ; the sample containing the brightest burst) among the FRB sources, J1846-0257 and J1854+0306 (Zhang et al., 2024a) among the pulsar sources, and J0534+2200 (Bera & Chengalur, 2019, the sample containing the largest number of bright bursts) and J0540-6919 (Geyer et al., 2021, ; the most extensive sample) as giant pulse comparison sources.
@@ 2.2 Data processing
where $D_{L}$ is the luminosity distance, $z$ is the redshift, and $F$ is the reported pulse fluence from the burst samples described above. The energy values for each source were then normalized to their brightest bursts, as our analysis focuses on comparing the high-energy tails.
Since only the cumulative energy distribution tail is available for the only nearby repeating FRB 20200120E, we used the power-law indices ($\alpha$) derived from the cumulative energy distribution in this study, which is given by:
where $N(\geq E)$ represents the predicted number of bursts with specific energy above $E$. In our distribution analysis, we adjusted the bins for each source to cover a similar range in log-10 space. To enable comparisons between different sources, we defined a density that is calculated as the number of bursts per bin divided by the total number of detected bursts.
For data of pulses from a single source (i.e., Figure 1), we considered only simple symmetric uncertainties in the counts, and fitted the power-law-like tails using the Python package curve_fit function from the SciPy library (Virtanen et al., 2020). The cutoff point for fitting was determined by identifying the location where the standard error was minimized. For data of pulses from different sources (i.e., Figure 2), we performed the fitting within a Bayesian framework using Markov Chain Monte Carlo (MCMC) sampling, via the emcee Python package (Foreman-Mackey et al., 2013). The fitting was conducted in log-log space using an asymmetric Gaussian likelihood.
@@ 3 Results
Figure 1: Cumulative specific energy distributions of high-energy bursts from FRB 20201124A and FRB 20121102A, single pulses from RRATs J1846-0257 and J1854+0306, and giant pulses from the Crab pulsar (J0534+2200) and J0540-6919. The energy values for each source are normalized to their respective brightest bursts. Each high-energy tail is fitted with a simple power law model. The fitting cutoff points are indicated by black dashed lines.
Figure 1 presents the cumulative specific energy distributions of high-energy bursts from the selected sources with relatively sufficient burst counts, most of which can be well-fitted by a simple power-law model. Specifically, the fitted power-law indices ($\alpha$) for FRB 20201124A (46 bursts [^3] For events detected simultaneously by multiple telescopes, we selected bursts from the observation with the largest bandwidth.), RRATs J1846-0257 (72 bright pulses) and J1854+0306 (325 bright pulses) are $0.67\pm 0.05$, $0.82\pm 0.03$ and $0.77\pm 0.01$, respectively, all of which are $\gtrsim-1$. In contrast, the indices for the Crab pulsar ($1.98\pm 0.07$, 1955 giant pulses) and J0540-6919 ($2.29\pm 0.1$, 865 bright single pulses) are significantly steeper, at $\lesssim-2$. The index for FRB 20121102 is $1.51\pm 0.04$ (1652 bursts), falling between the above two categories. Although the available data for bursts from FRB 20200120E and FRB 20220912A are insufficient to reprocess their high-energy distributions, the reported fitted power-law indices from cumulative distributions are $0.98\pm 0.06$ (Zhang et al., 2024b) for FRB 20200120E and range from 0.99-0.99 to 0.60-0.60 (Ould-Boukattine et al., 2024) for FRB 20220912A at around 1.4 GHz. Both values suggest relatively flat indices, consistent with the first group.